OpenStudy (anonymous):

The volume of a cylinder is 81 ft3. The volume of a cube is 42 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube?

OpenStudy (anonymous):

@hlee13 @helpme1.2

OpenStudy (anonymous):

@bcr1997

OpenStudy (anonymous):

not to good but lets see what I can do! shall WE

OpenStudy (anonymous):

do you know the volume of a cude and rectangle

OpenStudy (anonymous):

you first need to get the base area and height of the 2 figures originally given. then find the volume of the cone that fits exactly inside the cylinder. the cone would have the same base area as the cylinder as well as the same height, so just plug that into the area of cone formula: V = 1/3 * basearea * height the pyramid would likewise have the same base area and height plug those values in V = 1/3 * basearea * height and subtract those 2 values

OpenStudy (anonymous):

Volume = L x W x H L = Length W = Width H = Height Surface Area = 2(l x w) + 2 (L x h) + 2 (w x h)